### v6

@@ Appendix A Some details in the Proofs of Well-posedness and Mean-field limit
$dV^{i,N}_{t}=-\lambda\left(V^{i,N}_{t}-v_{\alpha}(\hat{\rho}_{t}^{N})\right)\,dt-\dfrac{1}{\epsilon}\nabla G\left(V^{i,N}_{t}\right)\,dt+\sigma\text{diag}\left(V^{i,N}_{t}-v_{\alpha}(\hat{\rho}_{t}^{N})\right)\,dB^{i,N}_{t},$
$V^{i,N}_{0}\sim\rho_{0},$
where i=1,,Ni=1,...,N. We can concantenate $\left\{V_{t}^{i,N}\right\}_{i=1}^{N}$ into one vector and put them in one equation. To be specific, we define
@@
where $\tilde{b}_{N}$ is some positive number that depends only on $\lambda,\sigma,d,\epsilon$ and $N$. Then we apply Theorem 3.1 in [17] to finish the proof.
A.2 Proof of Theorem 3.4
Below is Lemma 3.2 from [11]
@@
$+\mathbb{E}\left[\left(\int_{0}^{t}\big{\|}\nabla G(V_{\tau})-\nabla G(\hat{V}_{\tau})\big{\|}\,d\tau\right)^{2}\right]\mathbb{E}\left[\big{\|}\int_{0}^{t}\text{diag}\left(\left(V_{\tau}-u_{\tau}\right)-\left(\hat{V}_{\tau}-\hat{u}_{\tau}\right)\right)\,dB_{\tau}\big{\|}^{2}\right].$
@@
(52) $\mathbb{E}\left[(\int_{0}^{t}\big{\|}u_{\tau}-\hat{u}_{\tau}\big{\|}\,d\tau)^{2}\right]=\mathbb{E}\left[(\int_{0}^{t}\big{\|}v_{\alpha}(\rho_{\tau})-v_{\alpha}(\hat{\rho}_{\tau})\big{\|}\,d\tau)^{2}\right]\leq t\mathbb{E}\left[\int_{0}^{t}\big{\|}v_{\alpha}(\rho_{\tau})-v_{\alpha}(\hat{\rho}_{\tau})\big{\|}^{2}\,d\tau\right],$
@@
Then applying Gronwall’s inequality with the fact that $\mathbb{E}[\big{\|}Z_{0}\big{\|}^{2}]=0$ gives the uniqueness result.
A.3 Proof of Theorem 3.5
@@
$\mathbb{E}\big{\|}V^{i,N}_{t}\big{\|}^{2p}\lesssim$ $\mathbb{E}\big{\|}V_{0}^{i,N}\big{\|}^{2p}+\mathbb{E}(\int_{0}^{t}\big{\|}V^{i,N}_{\tau}\big{\|}\,d\tau)^{2p}+\mathbb{E}(\int_{0}^{t}\big{\|}v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\big{\|}\,d\tau)^{2p}$
$+\mathbb{E}\big{\|}\int_{0}^{t}\text{diag}\left(V^{i,N}_{\tau}\right)\,dB_{\tau}^{i}\big{\|}^{2p}+\mathbb{E}\big{\|}\int_{0}^{t}\text{diag}\left(v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\,dB_{\tau}^{i}\big{\|}^{2p}.$
@@
$\mathbb{E}(\int_{0}^{t}\big{\|}v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\big{\|}\,d\tau)^{2p}\leq t^{p}\cdot\mathbb{E}(\int_{0}^{t}\big{\|}v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\big{\|}^{2}\,d\tau)^{p}.$
@@
$\mathbb{E}\big{\|}\int_{0}^{t}\text{diag}\left(v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\,dB_{\tau}^{i}\big{\|}^{2p}=\mathbb{E}(\int_{0}^{t}\big{\|}v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\big{\|}^{2}\,d\tau)^{p}.$
@@
$\mathbb{E}\big{\|}V^{i,N}_{t}\big{\|}^{2p}\lesssim\mathbb{E}\big{\|}V_{0}^{i,N}\big{\|}^{2p}+\mathbb{E}(\int_{0}^{t}\big{\|}V^{i,N}_{\tau}\big{\|}^{2}\,d\tau)^{p}+\mathbb{E}(\int_{0}^{t}\big{\|}v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\big{\|}^{2}\,d\tau)^{p}.$
@@
$\mathbb{E}(\int_{0}^{t}\big{\|}V^{i,N}_{\tau}\big{\|}^{2}\,d\tau)^{p}\leq\mathbb{E}\int_{0}^{t}\big{\|}V^{i,N}_{\tau}\big{\|}^{2p}\,d\tau\text{ and }\mathbb{E}(\int_{0}^{t}\big{\|}v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\big{\|}^{2}\,d\tau)^{p}\leq\mathbb{E}\int_{0}^{t}\big{\|}v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\big{\|}^{2p}\,d\tau.$
@@
$\mathbb{E}\big{\|}V^{i,N}_{t}\big{\|}^{2p}\lesssim\mathbb{E}\big{\|}V_{0}^{i,N}\big{\|}^{2p}+\mathbb{E}\int_{0}^{t}\big{\|}V^{i,N}_{\tau}\big{\|}^{2p}\,d\tau+\mathbb{E}\int_{0}^{t}\big{\|}v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\big{\|}^{2p}\,d\tau.$
@@
(55) $\mathbb{E}\int\big{\|}v\big{\|}^{2p}\,d\hat{\rho}_{t}^{N}\lesssim\mathbb{E}\int\big{\|}v\big{\|}^{2p}\,d\hat{\rho}^{N}_{0}+\int_{0}^{t}(\mathbb{E}\int\big{\|}v\big{\|}^{2p}\,d\hat{\rho}_{\tau}^{N})\,d\tau+\int_{0}^{t}(\mathbb{E}\big{\|}v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\big{\|}^{2p})\,d\tau.$
@@
$\big{\|}v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\big{\|}^{2p}$ $=\big{\|}\int v\cdot\dfrac{\omega_{\alpha}(v)}{\big{\|}\omega_{\alpha}\big{\|}_{L^{1}(\tilde{\rho}_{\tau}^{N})}}\,d\hat{\rho}^{N}_{\tau}\big{\|}^{2p}$
@@
$\leq\left(\int\big{\|}v\big{\|}^{2}\cdot\dfrac{\omega_{\alpha}(v)}{\big{\|}\omega_{\alpha}\big{\|}_{L^{1}(\tilde{\rho}_{\tau}^{N})}}\,d\hat{\rho}^{N}_{\tau}\right)^{p}\leq\left(b_{1}+b_{2}\int\big{\|}v\big{\|}^{2}\,d\hat{\rho}^{N}_{\tau}\right)^{p}\lesssim 1+\int\big{\|}v\big{\|}^{2p}\,d\hat{\rho}^{N}_{\tau},$
@@
$\big{\|}v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\big{\|}^{2p}\lesssim 1+\int\big{\|}v\big{\|}^{2p}\,d\hat{\rho}^{N}_{\tau},$
@@
So $\mathbb{E}\big{\|}v_{\alpha}(\hat{\rho}^{N}_{t})\big{\|}^{2p}$ is bounded for $t\in[0,T]$ and the bound does not depend on $N$.
@@
$+\dfrac{1}{\epsilon}\int_{0}^{t}\left\langle\nabla G(v)\cdot\nabla\phi(v),\mu_{\tau}\right\rangle\,d\tau-\dfrac{\sigma^{2}}{2}\int_{0}^{t}\left\langle\sum_{k=1}^{d}\left(v-v_{\alpha}(\rho_{\tau})\right)_{k}^{2}\partial_{kk}\phi(v),\mu_{\tau}\right\rangle\,d\tau.$
@@
$F_{\phi}(\hat{\rho}_{t}^{N})=$ $\dfrac{1}{N}\sum_{i=1}^{N}\phi(V^{i,N}_{t})-\dfrac{1}{N}\sum_{i=1}^{N}\phi(V_{0}^{i,N})+\lambda\int_{0}^{t}\dfrac{1}{N}\sum_{i=1}^{N}\left(V^{i,N}_{\tau}-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\cdot\nabla\phi(V^{i,N}_{\tau})\,d\tau$
@@
$-\dfrac{\sigma^{2}}{2}\int_{0}^{t}\dfrac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{d}\left(V^{i,N}_{\tau}-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)_{kk}^{2}\partial_{kk}\phi(V^{i,N}_{\tau})\,d\tau.$
@@
$\phi(V^{i,N}_{t})-\phi(V_{0}^{i,N})=$ $-\lambda\int_{0}^{t}\left(V^{i,N}_{\tau}-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\cdot\nabla\phi(V^{i,N}_{\tau})\,d\tau-\dfrac{1}{\epsilon}\int_{0}^{t}\nabla G(V^{i,N}_{\tau})\cdot\nabla\phi(V^{i,N}_{\tau})\,d\tau$
$+\sigma\int_{0}^{t}\left(\nabla\phi(V^{i,N}_{\tau})\right)^{T}\left(\text{diag}\left(V^{i,N}_{\tau}-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\,dB^{i}_{\tau}\right)$
$+\dfrac{\sigma^{2}}{2}\int_{0}^{t}\sum_{k=1}^{d}\left(V^{i,N}_{\tau}-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)_{k}^{2}\partial_{kk}\phi(V^{i,N}_{\tau})\,d\tau.$
@@
$F_{\phi}(\hat{\rho}_{t}^{N})=\dfrac{\sigma}{N}\int_{0}^{t}\sum_{i=1}^{N}\left(\nabla\phi(V^{i,N}_{\tau})\right)^{T}\left(\text{diag}\left(V^{i,N}_{\tau}-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\,dB^{i}_{\tau}\right).$
@@
$\mathbb{E}\left[|F_{\phi}(\hat{\rho}_{t}^{N})|^{2}\right]$ $=\dfrac{\sigma^{2}}{N^{2}}\sum_{i=1}^{N}\mathbb{E}\left[\Big{|}\int_{0}^{t}\sum_{i=1}^{N}\left(\nabla\phi(V^{i,N}_{\tau})\right)^{T}\text{diag}\left(V^{i,N}_{\tau}-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\,dB^{i}_{\tau}\Big{|}\right]^{2}$
$=\dfrac{\sigma^{2}}{N^{2}}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{t}\sum_{i=1}^{N}\big{\|}\left(\nabla\phi(V^{i,N}_{\tau})\right)^{T}\text{diag}\left(V^{i,N}_{\tau}-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\big{\|}_{2}^{2}\,d\tau\right]$
$\leq\dfrac{\sigma^{2}}{N^{2}}\big{\|}\nabla\phi\big{\|}_{\infty}^{2}\sum_{i=1}^{N}\int_{0}^{t}\mathbb{E}\left[\big{\|}V^{1,N}_{\tau}-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\big{\|}_{2}^{2}\right]\,d\tau$
@@
where in the second equality, we used Itô’s isometry and in the third inequality, we used Lemma A.6. This completes the proof.
Now we recall the Aldous criteria ([5], Section 34.3), which can be used to prove the tightness of a sequence of distributions:
@@
(57) $V^{1,N}_{\beta+\delta}-V^{1,N}_{\beta}=$ $-\lambda\int_{\beta}^{\beta+\delta}\left(V_{\tau}^{1,N}-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\,d\tau+\sigma\int_{\beta}^{\beta+\delta}\text{diag}\left(V^{1,N}_{\tau}-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\,dB_{\tau}^{1}$
@@
(58) $\mathbb{E}\left[\big{\|}\lambda\int_{\beta}^{\beta+\delta}\left(V^{1,N}_{\tau}-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\,d\tau\big{\|}^{2}\right]\leq 2TK\lambda^{2}\delta$
@@
(59) $\mathbb{E}\left[\big{\|}\sigma\int_{\beta}^{\beta+\delta}\text{diag}\left(V^{1,N}-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\,dB_{\tau}^{1}\big{\|}^{2}\right]\leq\sigma^{2}\sqrt{8\delta TK}.$
@@
Choose $\delta_{0}$ small enough finishes the proof.
@@
  2. 2.
    $\mathbb{E}\left[\int_{0}^{t}\left\langle\left(v-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\cdot\nabla\phi(v),\hat{\rho}_{\tau}^{N}\right\rangle\,d\tau-\int_{0}^{t}\left\langle\left(v-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)\cdot\nabla\phi(v),\rho_{\tau}\right\rangle\,d\tau\right]$ converges to 0 as $N\rightarrow\infty$.
  3. 3.
    $\mathbb{E}\left[\int_{0}^{t}\left\langle\sum_{k=1}^{d}\left(v-v_{\alpha}\left(\hat{\rho}^{N}_{\tau}\right)\right)_{k}^{2}\partial_{kk}\phi(v),\hat{\rho}_{\tau}^{N}\right\rangle\,d\tau-\int_{0}^{t}\left\langle\sum_{k=1}^{d}\left(v-v_{\alpha}(\rho_{\tau})\right)_{k}^{2}\partial_{kk}\phi(v),\rho_{\tau}\right\rangle\,d\tau\right]$ converges to 0 as $N\rightarrow\infty$.
  4. 4.
@@
Then we utilize Lemma A.17 to establish that every convergent subsequence converges to the same limit: the unique solution to (8). Combining with Theorem A.12, we deduce that $\{\hat{\rho}_{t}^{N}\}_{N\in\mathbb{N}}$ converges and the limit is exactly the solution to (8).
A.4 Some auxiliary results used in the proof of Theorem 3.5
@@
This implies that $\int\psi\delta\rho_{t_{0}}=0$ for any chosen $\psi\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{d})$ and $t_{0}\in[0,T]$. Thus $\delta\rho_{t}=0.$ This proves the uniqueness.
@@
This completes the proof.

### v8

@@ Appendix A Some details in the Proofs of Well-posedness and Mean-field limit
$dV^{i,N}_{t}=$ $-\lambda\left(V^{i,N}_{t}-v_{\alpha}(\hat{\rho}_{t}^{N})\right)\,dt-\dfrac{1}{\epsilon}\nabla G\left(V^{i,N}_{t}\right)\,dt$
$+\sigma\text{diag}\left(V^{i,N}_{t}-v_{\alpha}(\hat{\rho}_{t}^{N})\right)\,dB^{i,N}_{t},$
with initial distribution $V^{i,N}_{0}\sim\rho_{0}$, where $i=1,\dots,N$. We can concantenate $\left\{V_{t}^{i,N}\right\}_{i=1}^{N}$ into one vector and put them in one equation. To be specific, we define
@@
where $\tilde{b}_{N}$ is some positive number that depends only on $\lambda,\sigma,d,\epsilon$ and $N$. Then we apply Theorem 3.1 in [18] to finish the proof. ∎
A.2. Proof of Theorem 3.3
Below is Lemma 3.2 from [12].
@@
$\quad\quad+$ $\mathbb{E}\left[\left(\int_{0}^{t}\big\|\nabla G(V_{\tau})-\nabla G(\hat{V}_{\tau})\big\|\,d\tau\right)^{2}\right]$
$\times\mathbb{E}\left[\big\|\int_{0}^{t}\text{diag}\left(\left(V_{\tau}-u_{\tau}\right)-\left(\hat{V}_{\tau}-\hat{u}_{\tau}\right)\right)\,dB_{\tau}\big\|^{2}\right].$
@@
(56) $\mathbb{E}\left[(\int_{0}^{t}\big\|u_{\tau}-\hat{u}_{\tau}\big\|\,d\tau)^{2}\right]$
$=\mathbb{E}\left[(\int_{0}^{t}\big\|v_{\alpha}(\rho_{\tau})-v_{\alpha}(\hat{\rho}_{\tau})\big\|\,d\tau)^{2}\right]\leq t\mathbb{E}\left[\int_{0}^{t}\big\|v_{\alpha}(\rho_{\tau})-v_{\alpha}(\hat{\rho}_{\tau})\big\|^{2}\,d\tau\right],$
@@
Then applying Gronwall’s inequality with the fact that $\mathbb{E}[\big\|Z_{0}\big\|^{2}]=0$ gives the uniqueness result. ∎
A.3. Proof of Theorem 3.4
@@
$\mathbb{E}\big\|V^{i,N}_{t}\big\|^{2p}\lesssim$ $\mathbb{E}\big\|V_{0}^{i,N}\big\|^{2p}+\mathbb{E}(\int_{0}^{t}\big\|V^{i,N}_{\tau}\big\|\,d\tau)^{2p}+\mathbb{E}(\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|\,d\tau)^{2p}$
$+\mathbb{E}\big\|\int_{0}^{t}\text{diag}\left(V^{i,N}_{\tau}\right)\,dB_{\tau}^{i}\big\|^{2p}+\mathbb{E}\big\|\int_{0}^{t}\text{diag}\left(v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB_{\tau}^{i}\big\|^{2p}.$
@@
$\mathbb{E}(\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|\,d\tau)^{2p}\leq t^{p}\cdot\mathbb{E}(\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2}\,d\tau)^{p}.$
@@
$\mathbb{E}\big\|\int_{0}^{t}\text{diag}\left(v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB_{\tau}^{i}\big\|^{2p}=\mathbb{E}(\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2}\,d\tau)^{p}.$
@@
$\mathbb{E}\big\|V^{i,N}_{t}\big\|^{2p}\lesssim\mathbb{E}\big\|V_{0}^{i,N}\big\|^{2p}+\mathbb{E}(\int_{0}^{t}\big\|V^{i,N}_{\tau}\big\|^{2}\,d\tau)^{p}+\mathbb{E}(\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2}\,d\tau)^{p}.$
@@
$\mathbb{E}(\int_{0}^{t}\big\|V^{i,N}_{\tau}\big\|^{2}\,d\tau)^{p}\leq\mathbb{E}\int_{0}^{t}\big\|V^{i,N}_{\tau}\big\|^{2p}\,d\tau$
and
$\mathbb{E}(\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2}\,d\tau)^{p}\leq\mathbb{E}\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2p}\,d\tau.$
@@
$\mathbb{E}\big\|V^{i,N}_{t}\big\|^{2p}\lesssim\mathbb{E}\big\|V_{0}^{i,N}\big\|^{2p}+\mathbb{E}\int_{0}^{t}\big\|V^{i,N}_{\tau}\big\|^{2p}\,d\tau+\mathbb{E}\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2p}\,d\tau.$
@@
(59) $\mathbb{E}\int\big\|v\big\|^{2p}\,d\hat{\rho}_{t}^{N}\lesssim\mathbb{E}\int\big\|v\big\|^{2p}\,d\hat{\rho}^{N}_{0}+\int_{0}^{t}(\mathbb{E}\int\big\|v\big\|^{2p}\,d\hat{\rho}_{\tau}^{N})\,d\tau+\int_{0}^{t}(\mathbb{E}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2p})\,d\tau.$
@@
$\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2p}$ $=\big\|\int v\cdot\dfrac{\omega_{\alpha}(v)}{\big\|\omega_{\alpha}\big\|_{L^{1}(\tilde{\rho}_{\tau}^{N})}}\,d\hat{\rho}^{N}_{\tau}\big\|^{2p}$
@@
$\leq\left(\int\big\|v\big\|^{2}\cdot\dfrac{\omega_{\alpha}(v)}{\big\|\omega_{\alpha}\big\|_{L^{1}(\tilde{\rho}_{\tau}^{N})}}\,d\hat{\rho}^{N}_{\tau}\right)^{p}$
$\leq\left(b_{1}+b_{2}\int\big\|v\big\|^{2}\,d\hat{\rho}^{N}_{\tau}\right)^{p}\lesssim 1+\int\big\|v\big\|^{2p}\,d\hat{\rho}^{N}_{\tau},$
@@
$\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2p}\lesssim 1+\int\big\|v\big\|^{2p}\,d\hat{\rho}^{N}_{\tau},$
@@
So $\mathbb{E}\big\|v_{\alpha}(\hat{\rho}^{N}_{t})\big\|^{2p}$ is bounded for $t\in[0,T]$ and the bound does not depend on $N$. ∎
@@
$+\dfrac{1}{\epsilon}\int_{0}^{t}\left\langle\nabla G(v)\cdot\nabla\phi(v),\mu_{\tau}\right\rangle\,d\tau$
$-\dfrac{\sigma^{2}}{2}\int_{0}^{t}\left\langle\sum_{k=1}^{d}\left(v-v_{\alpha}(\rho_{\tau})\right)_{k}^{2}\partial_{kk}\phi(v),\mu_{\tau}\right\rangle\,d\tau.$
@@
$F_{\phi}(\hat{\rho}_{t}^{N})=$ $\dfrac{1}{N}\sum_{i=1}^{N}\phi(V^{i,N}_{t})-\dfrac{1}{N}\sum_{i=1}^{N}\phi(V_{0}^{i,N})$
$+\lambda\int_{0}^{t}\dfrac{1}{N}\sum_{i=1}^{N}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\cdot\nabla\phi(V^{i,N}_{\tau})\,d\tau$
@@
$-\dfrac{\sigma^{2}}{2}\int_{0}^{t}\dfrac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{d}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)_{kk}^{2}\partial_{kk}\phi(V^{i,N}_{\tau})\,d\tau.$
@@
$\phi(V^{i,N}_{t})-\phi(V_{0}^{i,N})=$ $-\lambda\int_{0}^{t}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\cdot\nabla\phi(V^{i,N}_{\tau})\,d\tau$
$-\dfrac{1}{\epsilon}\int_{0}^{t}\nabla G(V^{i,N}_{\tau})\cdot\nabla\phi(V^{i,N}_{\tau})\,d\tau$
$+\sigma\int_{0}^{t}\left(\nabla\phi(V^{i,N}_{\tau})\right)^{T}\left(\text{diag}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB^{i}_{\tau}\right)$
$+\dfrac{\sigma^{2}}{2}\int_{0}^{t}\sum_{k=1}^{d}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)_{k}^{2}\partial_{kk}\phi(V^{i,N}_{\tau})\,d\tau.$
@@
$F_{\phi}(\hat{\rho}_{t}^{N})=\dfrac{\sigma}{N}\int_{0}^{t}\sum_{i=1}^{N}\left(\nabla\phi(V^{i,N}_{\tau})\right)^{T}\left(\text{diag}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB^{i}_{\tau}\right).$
@@
$\mathbb{E}\left[|F_{\phi}(\hat{\rho}_{t}^{N})|^{2}\right]$ $=\dfrac{\sigma^{2}}{N^{2}}\sum_{i=1}^{N}\mathbb{E}\left[\Big|\int_{0}^{t}\sum_{i=1}^{N}\left(\nabla\phi(V^{i,N}_{\tau})\right)^{T}\text{diag}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB^{i}_{\tau}\Big|\right]^{2}$
$=\dfrac{\sigma^{2}}{N^{2}}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{t}\sum_{i=1}^{N}\big\|\left(\nabla\phi(V^{i,N}_{\tau})\right)^{T}\text{diag}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\big\|_{2}^{2}\,d\tau\right]$
$\leq\dfrac{\sigma^{2}}{N^{2}}\big\|\nabla\phi\big\|_{\infty}^{2}\sum_{i=1}^{N}\int_{0}^{t}\mathbb{E}\left[\big\|V^{1,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|_{2}^{2}\right]\,d\tau$
@@
where in the second equality, we used Itô’s isometry and in the third inequality, we used Lemma A.4. This completes the proof. ∎
Now we recall the Aldous criteria (Section 34.3[5]), which can be used to prove the tightness of a sequence of distributions:
@@
(61) $V^{1,N}_{\beta+\delta}-V^{1,N}_{\beta}=$ $-\lambda\int_{\beta}^{\beta+\delta}\left(V_{\tau}^{1,N}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,d\tau+\sigma\int_{\beta}^{\beta+\delta}\text{diag}\left(V^{1,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB_{\tau}^{1}$
@@
(62) $\mathbb{E}\left[\big\|\lambda\int_{\beta}^{\beta+\delta}\left(V^{1,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,d\tau\big\|^{2}\right]\leq 2TK\lambda^{2}\delta$
@@
(63) $\mathbb{E}\left[\big\|\sigma\int_{\beta}^{\beta+\delta}\text{diag}\left(V^{1,N}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB_{\tau}^{1}\big\|^{2}\right]\leq\sigma^{2}\sqrt{8\delta TK}.$
@@
Choose $\delta_{0}$ small enough finishes the proof. ∎
@@
  2. (2)
    $\mathbb{E}\left[\int_{0}^{t}\left\langle\left(v-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\cdot\nabla\phi(v),\hat{\rho}_{\tau}^{N}\right\rangle\,d\tau-\int_{0}^{t}\left\langle\left(v-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\cdot\nabla\phi(v),\rho_{\tau}\right\rangle\,d\tau\right]$ converges to 0 as $N\rightarrow\infty$.
  3. (3)
    The quantity
    $\mathbb{E}\Bigg[$ $\int_{0}^{t}\left\langle\sum_{k=1}^{d}\left(v-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)_{k}^{2}\partial_{kk}\phi(v),\hat{\rho}_{\tau}^{N}\right\rangle\,d\tau$
    $-\int_{0}^{t}\left\langle\sum_{k=1}^{d}\left(v-v_{\alpha}(\rho_{\tau})\right)_{k}^{2}\partial_{kk}\phi(v),\rho_{\tau}\right\rangle\,d\tau\Bigg]$
    converges to 0 as $N\rightarrow\infty$.
  4. (4)
@@
Then we utilize Lemma A.10 to establish that every convergent subsequence converges to the same limit: the unique solution to (8). Combining with Theorem A.8, we deduce that $\{\hat{\rho}_{t}^{N}\}_{N\in\mathbb{N}}$ converges and the limit is exactly the solution to (8). ∎
A.4. Some auxiliary results used in the proof of Theorem 3.4
@@
This implies that $\int\psi\delta\rho_{t_{0}}=0$ for any chosen $\psi\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{d})$ and $t_{0}\in[0,T]$. Thus $\delta\rho_{t}=0.$ This proves the uniqueness. ∎
@@
This completes the proof. ∎